OpenStudy (anonymous):

It’s About Time, in Langley, British Columbia, is Canada’s largest custom clock manufacturer. They have a grandfather clock that, on the hours, chimes the number of times that corresponds to the time of day. For example, at 4:00 p.m., it chimes 4 times. How many times does the clock chime in a 24-h period?

OpenStudy (anonymous):

Please use the concept of series and sequence here!

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Not at all. Can't figure out this sentence: They have a grandfather clock that, on the hours, chimes the number of times that corresponds to the time of day.

OpenStudy (skullpatrol):

At one o'clock it chimes 1 time At two o'clock it chimes 2 times At three o'clock it chimes 3 times . . At twelve o'clock it chimes 12 times

OpenStudy (anonymous):

So, in 24 hours, it'll be t24 = 1 + (24-1)1 = 1+23? = 24 but the answer's 156.

OpenStudy (skullpatrol):

$$\Huge 1 + 2 + 3 +...+11 + 12=12hours$$

OpenStudy (anonymous):

Shouldn't the series go up to 24 hours?

OpenStudy (skullpatrol):

That^ is the number of times in 12hours, right?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

You need to multiply that sum 1+2+3+...+11+12 by 2 to get the number of times the clock chimes in 24 hours, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but how do you put it in an equation and what is the question really looking for? That's what I can't understand. Shouldn't the clock chime 24 times for 24 hours?

OpenStudy (skullpatrol):

Why would " the clock chime 24 times for 24 hours?"

OpenStudy (anonymous):

I have no idea. I still don't understand the question. This is all I can deduce from the question and it's wrong!

OpenStudy (skullpatrol):

Ok, let us start at 12:01pm

OpenStudy (skullpatrol):

0 chimes right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

at 1, it'll chime once?

OpenStudy (skullpatrol):

yes, record 1 chime

OpenStudy (anonymous):

at 2, it'll chime twice.

OpenStudy (anonymous):

I understood that pattern.

OpenStudy (skullpatrol):

yes, record 2 and now ADD 1 + 2 = 3chimes so far

OpenStudy (anonymous):

for the first two terms.

OpenStudy (skullpatrol):

yes, keep going...

OpenStudy (anonymous):

4+5+6....

OpenStudy (skullpatrol):

yes, add them all up

OpenStudy (anonymous):

up to 6?

OpenStudy (skullpatrol):

no, up to midnight

OpenStudy (anonymous):

so, 12

OpenStudy (skullpatrol):

add 1+2+3+4+5+6+7+8+9+10+11+12

OpenStudy (anonymous):

it's 12

OpenStudy (skullpatrol):

that will take you to midnight

OpenStudy (anonymous):

and then?

OpenStudy (skullpatrol):

does 1+2+3+4+5+6+7+8+9+10+11+12=12?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

how?

OpenStudy (anonymous):

sorry it's 78.

OpenStudy (skullpatrol):

thank you

OpenStudy (skullpatrol):

now that is from 12:01pm to midnight which is a 12 hour period, the question asks for a 24 hour period, right?

OpenStudy (anonymous):

instead of manually adding them up, will the Sn= n/2 [2t1 + (n-1) d] formula work?

OpenStudy (skullpatrol):

Answer my question please :)

OpenStudy (anonymous):

yes. i got 300

OpenStudy (skullpatrol):

78 chimes times 2 =?= 300

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

no wait!

OpenStudy (anonymous):

haha i added them up myself and i got 300. what's the difference between adding them and multiplying the numbers. Shouldn't we be getting the same answers? > 300?

OpenStudy (skullpatrol):

nope